OpenStudy (anonymous):

Divide -3x3 - 2x2 - x - 2 by x – 2.

OpenStudy (anonymous):

A. -3x2 + 4x + 15 B. -3x2 + 4x + 15, R 32 C. -3x2 - 8x - 17 D. -3x2 - 8x - 17, R –36

OpenStudy (anonymous):

@neer2890

OpenStudy (neer2890):

you want shortcut or a proper method...?

OpenStudy (anonymous):

the easiest way to understand lol

OpenStudy (anonymous):

are you still there

OpenStudy (vishweshshrimali5):

One method is to use remainder theorem: It says that the remainder obtained on dividing some polynomial P (x) by (x-a) is P(a) i.e. in polynomial P(x) put a in place of x and then solve it.

OpenStudy (vishweshshrimali5):

For example, here in this case: P(x) = -3x^3 - 2x^2 - x - 2 we are dividing it by (x-2) , so, the remainder would be: P(2) Now, P(2) = -3*(2)^3 - 2(2)^2 - 2 -2

OpenStudy (vishweshshrimali5):

Can you solve P(2) ?

OpenStudy (anonymous):

-36

OpenStudy (vishweshshrimali5):

Good See: \[\large{P(2) = -3*(2)^3 - 2(2)^2 - 2 - 2}\] \[\large{\implies P(2) = -3*8 - 2*4 - 2 - 2}\] \[\large{\implies P(2) = -24 - 8 - 2 - 2}\] \[\large{\implies P(2) = -36}\]

OpenStudy (vishweshshrimali5):

So, remainder = -36

OpenStudy (vishweshshrimali5):

Now, if there would have been more than 1 option with remainder -36 , then, we would have applied division theorem: \[\large{\text{Dividend} = \text{Quotient}\times \text{Divisor} + \text{Remainder}}\]

OpenStudy (vishweshshrimali5):

Now, you have : \[\large{\text{Dividend} = -3x^3 -2x^2 - x -2}\] \[\large{\text{Divisor} = x-2}\] \[\large{\text{Remainder} = -36}\] \[\large{\text{Quotient} = ?}\]

OpenStudy (vishweshshrimali5):

So, can you calculate the quotient ?

OpenStudy (anonymous):

i don't know how to do that

OpenStudy (vishweshshrimali5):

All right lets see: \[\large{\text{Dividend} = \text{Divisor}\times \text{Quotient} + \text{Remainder}}\] \[\large{\implies\text{Dividend} - \text{Remainder} = \text{Divisor}\times \text{Quotient}}\] \[\large{\implies \text{Quotient} = \cfrac{\text{Dividend}-\text{Remainder}}{\text{Divisor}}}\]

OpenStudy (vishweshshrimali5):

Now, I am going to put the values

OpenStudy (vishweshshrimali5):

\[\large{\implies \text{Quotient} = \cfrac{(-3x^3 -2x^2 -x - 2) - (-36)}{(x-2)}}\] \[\large{\implies \text{Quotient} = \cfrac{-3x^3 - 2x^2 - x - 2 + 36}{x-2}}\] \[\large{\implies \text{Quotient} = \cfrac{-3x^3 - 2x^2 - x + 34}{x-2}}\]

OpenStudy (vishweshshrimali5):



OpenStudy (vishweshshrimali5):

Thus, \[\large{\text{Quotient} = -3x^2 - 8x - 17}\]

OpenStudy (vishweshshrimali5):

Thus, D is the correct option. Now do you get it ?

OpenStudy (anonymous):

ok i see so it is D, thank you so much for helping but i have another question can you help me on that one too